Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 65     Entry time: 02 Mar 2023, 23:13
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 202 pulses, X-ray intensity had slope over train, 30 % difference between first and last pulse in the train
  • Sometime strange intensity slope was observed (screen shot was attached)

Optical laser delivery

  • 800 nm, 68 pulses, 188 kHz, Treacy compressor issue, lost 3 hours in morning
  • ns laser for illumination

Achievements/Observations

  • Nanofocusing setup with pump-probe experiment has started
  • Collected TR-SFX data

Issues

  • Treacy compressor issue, lost 3 hours in morning
  •  X-ray intensity had slope over train, 30 % difference between first and last pulse in the train
  • Sometime strange intensity slope was observed  (screen shot was attached)
  • Run_Controller stopped without any action from us
Attachment 1: Screenshot_from_2023-03-02_21-53-17.png  62 kB  | Hide | Hide all
Screenshot_from_2023-03-02_21-53-17.png
ELOG V3.1.4-7c3fd00